OpenStudy (anonymous):

The excursion boat on the river takes 2½ hours to make the trip to a point 12 miles upstream and to return. If the rate at which the boat travels in still water is 5 times the rate of the river current, what is the rate of the current?

OpenStudy (anonymous):

assume the river current as V,so the boat's velocity is 5*V.The boat travels against the upstream so it have a relative velocity 5*V - V = 4*V The speed is given 4V = 12/2.5 (miles/h), V = 1.2 miles/h